OpenStudy (anonymous):

Convert the following binary number to decimal and hexadecimal assuming 16-bits for each number and two's complement representation for negative numbers: 11101001001111002

OpenStudy (anonymous):

researching what i need to do here. if anyone wants to help that would be greatly appreciated! :)

OpenStudy (anonymous):

@jim_thompson5910 :)

OpenStudy (freckles):

what is the binary number you talk about

OpenStudy (freckles):

or am I suppose to imagine that 2 at the end there is not there

OpenStudy (freckles):

binary numbers are numbers made up of only one's and zero's

OpenStudy (anonymous):

that two is representing base two

OpenStudy (freckles):

example of converting binary to base 10: \[11101=1 \cdot 2^4+1 \cdot 2^3+1 \cdot 2^2+0 \cdot 2^1+1 \cdot 2^{0}=8+4+1=29 \\ \\ \text{ so this means } 11101_2=29_{10} \\ \text{ hmm..,. maybe that 2 at the end is a subscript you have above }\] anyways I usually convert to hexadecimal after converting to base 10... \[29=1 \cdot 16^1+13 \cdot 16^0 \\ \text{ we need a letter for 13 } \\ 10=a\\ 11=b\\12=c\\ 13=d \\ 14=e \\ 15=f \\ \text{ so we have } 29_{10}=1d_{16}\]

OpenStudy (anonymous):

i'm reading through this example. hold on

OpenStudy (freckles):

there is an easier way for bigger numbers for converting base 10 to whatever base...

OpenStudy (anonymous):

@freckles cool, so i raise 2 to power how many digets are behind the number i'm looking at and add. what is with the 16bit and two's compliment about?

OpenStudy (freckles):

but let me choose a long example because it might be hard to see a pattern we just two steps... say we wanted to write 29 in base 2... we already know what it should look like from doing the above example: